Fermi National Laboratory


Accelerator Update

Monday August 2
The day shift began with the Tevatron (TeV) in a 4x0 study store, with Operations monitoring the Recycler stash, but with no beam to Switchyard (SY) due to the upcoming shot setup.

Operations put the TeV into shot setup at 9 AM.

At 10:57 AM, a Booster West Gallery power supply (WAPS) tripped off due to a 480-volt breaker trip. The Duty Electrician and the High Voltage Coordinator were contacted. Repairs began at 11:23 AM. A Booster RF station (BRF15) went down with this power supply trip and never came back. It was bypassed from the RF system.

Around 11:20 AM, a Recycler power supply (R:V501) tripped off and caused the loss of stashed antiprotons. This power supply was left alone since any action would have compromised the remaining beam.

Operations established store 3699 at 11:42 AM with an initial luminosity of 66.93E30.

Operations established beam to the Meson MCenter experiment at 12:37 PM.

Operations made MiniBooNE ready for access at 1:30 PM.

The Antiproton Source (Pbar) began stacking at 1:34 PM.

A Main Injector (MI) expert ran beam to the MI dump at 2:10 PM for a tune up and then began a slipstacking study. The study ended at 4:27 PM.

Operations began sending beam to MiniBooNE at 6:19 PM. The Horn was off due to a ground fault.

Operations reported at 10:23 PM that they had been working on a Mixed Mode Stacking and possible Time Line Generator (TLG) problems. At 11:27 PM, operators changed out a control card for two power supplies; this may have helped a little.

Tuesday August 3
The midnight shift began with Operations monitoring store 3699 and the stash, with Pbar stacking, and with MiniBooNE, MCenter, and MTest taking beam.

At 1:17 AM, BRF13 tripped off. It reset okay.

At 4:50 AM, a Linac RF station (LRF1) tripped off. It reset okay.

At 4:57 AM, a TeV RF station (TRF7) tripped off. It will be reset after the store.

At 7:34 AM, LRF4 tripped off. It reset itself okay.

At 10:16 AM, BRF13 tripped off. It reset okay.

MiniBooNE beam was disabled at 10:39 AM so that experts could conduct an hour of Horn tests. (These tests later revealed that MiniBooNE could not run the Horn in this condition.)

Operations terminated store 3699 at noon.

At 1:38 PM, the TeV suffered a sector D1 quench.

At 1:44 PM, Operations reported that 120mA of the stack was lost. Investigations continued, but it looked like the stack had been over cooled.

Pbar resumed stacking at 1:58 PM.

Operations prepared the Collider Detector at Fermilab (CDF) and D-Zero (D0) for access at 2:13 PM.

Operations resumed sending beam to MCenter and MTest at 2:14 PM.

At 2:19 PM, a Linac Klystron RF station (KRF4) tripped off due to a coil failure. Linac technicians immediately began work on this station. This halted all down stream beam.

Techs completed the KRF4 work and Pbar resumed stacking at 3:04 PM.

Both CDF and D0 were out of access by 3:52 PM.

Operations put the TeV through a dry squeeze at 4:32 PM.

A TeV expert began a study around 5:40 PM.

At 7:33 PM, BRF5 tripped off a number of times over a fifteen-minute period.

At 8:17 PM, Operations reported that beam was not making it into the Accumulator.

At 8:52 PM, Operations reported that the TeV sector C4 readbacks came back and that the vacuum in that area had gone bad. Operations contacted experts who said to give it thirty minutes.

Operations reported at 9:26 PM that the Accumulator beam problem was due to a kicker (D:EKIK) that had tripped off. It wouldn't reset. Operators called in experts.

A lightning storm rolled through the area at 9:43 PM. The MI abort kickers tripped off. They reset okay.

Operations prepared the TeV for a sector C4 access to check for vacuum problems.

At 10 PM, MI suffered a ground fault due to a lightning strike. The upper VCB tripped at MI60 and it required many resets before it finally took. MiniBooNE also tripped off, but it reset right away.

At 10:09 PM, a TeV sector E2 wet engine and cryo compressor tripped off. Cryo system techs at CHL reset them.

At 10:40 PM, a Meson power supply (MW1W), which was clamped off for an experiment access, wouldn't come back on. Operators investigated and had beam back by 11:22 PM. However there still seemed to be a LCW problem. Operators called in an expert.

CHL reported at 10:59 PM that they lost control of their tank farm valves due to storm damage. The said they had started repairs.

Wednesday August 4
The midnight shift began with Operations monitoring the stash, with the TeV in access to fix vacuum problems, with techs working on the Meson LCW system, with an expert repairing a Debuncher kicker, and with MiniBooNE taking beam.

MCenter experimenters said that they were done with beam at 1 AM.

A vacuum tech in the TeV reported at 1:33 AM that there was a stuck Kautzky valve at D1. Operators informed the cryo system techs. Operators accessed the TeV at 3 AM and thawed the Kautzky valve body and control pressure line.

Operations reported at 4 AM that TeV turn on was in progress. Unfortunately, at 4:34 AM a power glitch affected Cryo, Linac, Booster, and the TeV RF. At 4:42 AM, the Duty Electrician discovered two phases of feeder 47 had tripped. He contacted the High Voltage Coordinator. Several cryo compressors tripped of around the ring.

TeV and Booster RF were back by 4:41 AM.

Operations reestablished beam to MCenter at 5:14 AM.

At 5:43 AM, BRF5 tripped off for the first of six times in five minutes. It reset okay.

Operations reestablished beam to MTest at 7 AM.

At 7:50 AM, the High Voltage Coordinator discovered a TeV sector E2 transformer-to-bus connection blown apart. Operations contacted EE support techs to begin repairs.

Plans for Wednesday
The plans for today are to get the TeV back into operating condition as soon as possible. The Run Coordinator will allow a TLG reboot to occur during the TeV work. A two to three-hour access is planned for tomorrow morning, 8/5/04, at 8 AM.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 8/4/2004   email Fermilab